# Unified Modeling and Structural Optimization of Multi-magnet Embedded Soft Continuum Robots for Enhanced Kinematic Performances

### v1

@@ III-B Optimization Methodology
The position and orientation of embedded magnets determine the magnetic response deformation of MeSCRs under the same magnetic field. For instance, embedded magnets with variable positions have been employed to expand the MeSCR’s workspace [8]. In the PRB model, the length of embedded magnets can be defined by the Voronoi length as $L_{k}\triangleq\sum_{i\leq k}\ell_{i},\forall k\in\mathbb{K}$. From (10), it follows that the flexible segment beyond the last embedded magnet is unactuated. Consequently, it does not affect the workspace volume of the MeSCR, nor any performance metric defined using a left-invariant Riemannian metric. Therefore, one of the magnets is fixed at the distal end of the continuum ($L_{k_{N_{m}-1}}=L$), leaving $N_{m}-1$ decision variables. Additionally, let the flexible segments between adjacent magnets be uniformly distributed and have appropriate spacing to mitigate magnetic coupling effects. Based on the previous assumption regarding the overall reversal of the magnetic field direction, any configuration of axially placed magnets has a corresponding symmetrical configuration, resulting in a total of $2^{N_{m}-1}$ possible combinations.
Expanding the workspace volume is a key objective in the optimal design of MeSCRs [43]. However, comparing workspace volumes across MeSCRs with different embedded magnet orientations is of limited significance, unless the focus is explicitly on spatial coverage by the distal end. As shown in Fig. 5, we compared the normalized half-workspace areas of the four MeSCRs under equidistant configurations. The results indicate that ② achieves the largest normalized workspace area, whereas ③ has the smallest. Nevertheless, despite its smaller workspace area, ① may still reach regions that are inaccessible to ②. Therefore, our focus shifts to the global indices of MeSCRs with fixed magnet orientations within their respective workspaces.
Recall that ${}_{s}^{b}\mathbf{H}(\boldsymbol{\theta}):\mathbb{T}\to\mathrm{SE}(3)$ is the forward kinematic map of the MeSCR. Let $z({}_{s}^{b}\mathbf{H}(\boldsymbol{\theta});L_{k}):\mathbb{T}\to\mathbb{R}$ be an integrable scalar functional defined on $\mathbb{T}$, representing a local performance density of ${}_{s}^{b}\mathbf{H}(\boldsymbol{\theta})$ with left-invariant Riemannian metric. This local measure can be extended to a global measure $\mathcal{Z}(L_{k})$ by integrating $z$ over $\mathbb{T}$, yielding the following optimization problem on $L_{k}$:
$\min_{L_{k}}$ $\ \mathcal{Z}(L_{k})\triangleq\int_{\mathbb{T}}z({}_{s}^{b}\mathbf{H}(\boldsymbol{\theta});L_{k})\ \Omega_{\mathbb{T}}$ (18)
[…]
where $\Omega_{\mathbb{T}}$ represents the Riemannian volume form on $\mathbb{T}$ induced from the Euclidean metric and $\mathcal{C}(L_{k})$ denotes linear constraints on $L_{k}$. The equilibrium constraint restricts feasible configurations to the solution set $\mathbb{T}^{*}\triangleq\{\boldsymbol{\theta}\in\mathbb{T}|\nabla_{\boldsymbol{\theta}}\mathcal{E}(\boldsymbol{\theta},\mathbf{b})=\mathbf{0}\}$. If the actuation manifold $\mathbb{B}$ is arbitrary, the set $\mathbb{T}^{*}$ may not admit an explicit parametrization. Nevertheless, under mild regularity conditions on $\mathcal{E}$, and by Sard’s theorem, the solution set $\mathbb{T}^{*}\subset\mathbb{T}$ is countably $d$-rectifiable, with $d=3N-\operatorname{rk}(\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}})$. Hence, it admits a well-defined $\mathcal{H}^{d}$-measure and supports geometric integration. Consequently, the constrained integral over $\mathbb{T}$ may be rewritten as an iterated integral over $\mathbb{B}$ and the corresponding solution manifold:
$\mathcal{Z}(L_{k})=\int_{\mathbb{B}}\left(\int_{\mathbb{T}^{*}}z(_{s}^{b}\mathbf{H}(\boldsymbol{\theta});L_{k})\mathrm{d}\mathcal{H}^{d}(\boldsymbol{\theta})\right)\Omega_{\mathbb{B}}.$
Now, suppose the external magnetic field $\mathbf{b}$ is spatially uniform with $\mathbf{b}=\mathbf{U}_{u}\mathbf{b}_{u}$ and satisfies the condition in Theorem 1. The manifold $\mathbb{B}$ is now an affine subspace diffeomorphic to $\mathbb{R}^{3}$. By Theorem 1, the stiffness matrix $\mathbf{S}$ is nonsingular in a neighborhood of each $\mathbf{b}_{u}\in\mathbb{B}$. By Theorem 3, the Jacobian $\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}}=\mathbf{S}^{-1}\mathbf{M}$ satisfies $\operatorname{rk}(\mathbf{S}^{-1}\mathbf{M})=2N_{m}$, which ensures that $\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}_{u}}=\mathbf{S}^{-1}\mathbf{M}\mathbf{U_{u}}$ is of full rank provided $N_{m}\geq 2$. In this regime, the equilibrium constraint gives rise to a smooth immersion $\boldsymbol{\theta}^{*}:\mathbb{B}\to\mathbb{T}$. The inner integral reduces to an evaluation at $\boldsymbol{\theta}^{*}(\mathbf{b})$, and using the area formula [44] yields:
$\int_{\mathbb{T}^{*}}z(_{s}^{b}\mathbf{H}(\boldsymbol{\theta});L_{k})\mathrm{d}\mathcal{H}^{d}(\boldsymbol{\theta})=z(_{s}^{b}\mathbf{H}(\boldsymbol{\theta});L_{k})\mathcal{J}_{\boldsymbol{\theta}}(\mathbf{b}),$
where $\mathcal{J}_{\boldsymbol{\theta}}(\mathbf{b})$ denotes the Jacobian of the immersion $\boldsymbol{\theta}^{*}(\mathbf{b})$ and is given by
$\mathcal{J}_{\boldsymbol{\theta}}=\left[\det\left(\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}_{u}}^{\top}\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}_{u}}\right)\ \right]^{\frac{1}{2}}\triangleq\left[\mbox{det}\left(\mathbf{G}_{\boldsymbol{\theta}}\right)\right]^{\frac{1}{2}}.$
where $\mathbf{G}_{\boldsymbol{\theta}}$ denotes the Gram matrix of $\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}_{u}}$. Substituting back, the global optimization problem (18) takes the following unconstrained reformulation:
$\min_{L_{k}}$ $\ \mathcal{Z}(L_{k})=\int_{\mathbb{B}}z({}_{s}^{b}\mathbf{H}(\boldsymbol{\theta}^{*}(\mathbf{b}));L_{k})\mathcal{J}_{\boldsymbol{\theta}}\ \Omega_{\mathbb{B}}.$ (19)
The optimality is ensured by satisfying the following first-order necessary condition:
$\nabla_{L_{k}}\mathcal{Z}$ $=\int_{\mathbb{B}}\left(z\nabla_{L_{k}}\mathcal{J}_{\boldsymbol{\theta}}+\mathcal{J}_{\boldsymbol{\theta}}\nabla_{L_{k}}z\right)\ \Omega_{\mathbb{B}}$
$=\int_{\mathbb{B}}\left[\frac{1}{2}z\mathcal{J}_{\boldsymbol{\theta}}\operatorname{tr}\left(\mathbf{G}_{\boldsymbol{\theta}}^{-1}\nabla_{L_{k}}\mathbf{G}_{\boldsymbol{\theta}}\right)+\mathcal{J}_{\boldsymbol{\theta}}\nabla_{L_{k}}z\right]\ \Omega_{\mathbb{B}}$
$=0,\ \forall k\in\mathbb{K}/\{N_{m}-1\}.$
Under regularity conditions, this implies the strong-form pointwise optimality condition:
$\left(\frac{1}{2}z\operatorname{tr}\left(\mathbf{G}_{\boldsymbol{\theta}}^{-1}\nabla_{L_{k}}\mathbf{G}_{\boldsymbol{\theta}}\right)+\nabla_{L_{k}}z\right)\mathcal{J}_{\boldsymbol{\theta}}=0,\ \forall\mathbf{b}_{u}\in\mathbb{B}.$ (20)
@@
Observed that the condition trivially holds when $\mathcal{J}_{\boldsymbol{\theta}}=0$. However, the collapse of immersion Jacobian results in a degenerate configuration. That is, the embedded magnets may have merged or be unable to generate torques. Since $\mathcal{J}_{\boldsymbol{\theta}}$ contains complex Hessian terms, we first derive a tractable approximation to facilitate further analysis.
@@
$\mathcal{J}_{\boldsymbol{\theta}}^{2}$ $=\det\left(\mathbf{U}_{u}^{\top}\mathbf{M}^{\top}\boldsymbol{\Lambda}^{-2}\mathbf{M}\mathbf{U}_{u}\right)+\epsilon$ (21)
$\triangleq\tilde{\mathcal{J}}_{\boldsymbol{\theta}}^{2}+\epsilon$
@@
The proof is deferred to the supplementary material. ∎
It shows that $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}$ provides a close approximation of $\mathcal{J}_{\boldsymbol{\theta}}$ in the weak-field regime, with the residual term converging to zero as the magnetic field strength decreases. We can now turn to analyze the approximation of the objective $\tilde{\mathcal{Z}}(L_{k})\triangleq\int_{\mathbb{B}}z\ \tilde{\mathcal{J}}_{\boldsymbol{\theta}}\ \Omega_{\mathbb{B}}$. Let $\mathbf{M}(\boldsymbol{\theta})$ in (12) be represented as $\mathbf{M}=\mathbf{D}_{k_{m}-1}^{\top}\mathbf{U}_{m}\mathbf{D}_{m}$, we notice
@@
with $\Lambda_{k_{i}}^{m}=\sum_{j=0}^{k_{i}}\frac{\ell_{j}^{2}}{E_{j}^{2}I_{j}^{2}}\approx\frac{1}{k_{0}E^{2}I^{2}}\sum_{j=0}^{i}(L_{k_{j}}-L_{k_{j-1}})^{2}$ as a function of $L_{k}$. The following lemma shows the relation between the singularity of $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}$ and the configuration of $L_{k}$:
@@
[…]
$\left(\mathbf{v}^{\top}\otimes\mathbf{I}_{3}\right)\mathbf{U}_{u}\mathbf{D}_{m}=\sum_{i=N_{k}+1}^{N_{m}-1}\left[\mathbf{m}_{k_{i}}\right]_{\times}=\left[\sum_{i=N_{k}+1}^{N_{m}-1}\mathbf{m}_{k_{i}}\right]_{\times}.$
Consequently, the rank of the expression for $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}$ shows that
$\operatorname{rk}\!\left(\mathbf{U}_{u}^{\top}\mathbf{D}_{m}^{\top}\!\left(\boldsymbol{\Lambda}_{m}\!\otimes\mathbf{I}_{3}\right)\mathbf{U}_{u}\mathbf{D}_{m}\right)\!$
$=\operatorname{rk}\!\left(\!\left(\mathbf{v}^{\top}\!\!\otimes\mathbf{I}_{3}\right)\!\mathbf{U}_{u}\mathbf{D}_{m}\right)=\operatorname{rk}\left(\left[\sum_{i=N_{k}+1}^{N_{m}-1}\mathbf{m}_{k_{i}}\right]_{\times}\right)\leq 2.$
That is, the actuation mapping becomes deficient in rank, and we conclude that $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}=0$. ∎
@@
If the performance index satisfies $z(\mathbf{b}_{u};L_{k})>0$, and the structural parameter $L_{k}$ satisfies the boundary configuration condition in Lemma 4, then the objective $\tilde{\mathcal{Z}}(L_{k})$ attains its minimum at this configuration.
@@
Since $z>0$ and $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}\geq 0$, it follows that $\tilde{\mathcal{Z}}(L_{k})\geq 0$ for any admissible $L_{k}$. When $L_{k}$ satisfies the boundary configuration condition in Lemma 4, we have $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}(\mathbf{b})\equiv 0$ over $\mathbb{B}$, and thus $\tilde{\mathcal{Z}}(L_{k})=0$. ∎
@@
Due to the Jacobian singularity at boundary configurations, the value of the objective becomes insensitive to the gradient of the performance index, thereby weakening the meaningfulness of the solution. However, the minimum spacing constraint prevents the structural parameter $L_{k}$ from reaching such degenerate configurations. Since $\mathcal{J}_{\boldsymbol{\theta}}$ remains continuous on $L_{k}$, it stays small in the vicinity of the boundary, making nearby configurations highly competitive in minimizing the objective. In this case, the performance gradient becomes the key factor in selecting specific configurations. One can also reduce the optimization bias by considering the objective as $\mathcal{Z}/\mathcal{V}$, where $\mathcal{V}\triangleq\int_{\mathbb{B}}\mathcal{J}_{\boldsymbol{\theta}}\Omega_{\mathbb{B}}$ is the total induced volume. This normalized objective reflects the average performance on the configuration space and is insensitive to Jacobian singularity.
Now we focus on the nontrivial solution of (20). If $z$ is a non-zero performance index, one can further deduce that
$\operatorname{tr}(\mathbf{G}_{\boldsymbol{\theta}}^{-1}\nabla_{L_{k}}\mathbf{G}_{\boldsymbol{\theta}})=\sum_{i}\frac{\nabla_{L_{k}}\lambda_{i}(\mathbf{G}_{\boldsymbol{\theta}})}{\lambda_{i}(\mathbf{G}_{\boldsymbol{\theta}})}=-2\frac{\nabla_{L_{k}}z}{z}.$ (22)
Geometrically, this condition implies that optimizing the local performance requires structurally modulating the spectrum of the configuration space metric to counteract its distortion. From Lemma 3, we have the following approximation:
$=\sum_{ij}\Lambda_{ij}^{m}\left(\mathbf{m}_{k_{i}}\mathbf{m}_{k_{j}}^{\top}-\left(\mathbf{m}_{k_{i}}^{\top}\mathbf{m}_{k_{j}}\right)\mathbf{I}_{3}\right)$
in which $\mathbf{m}_{k}=M(\cos(\theta_{0}^{k})\bar{\mathbf{t}}-\sin(\theta_{0}^{k})\mathbf{v})$. While the spectrum is analytically elusive for general $\mathbf{v}\in\mathbb{S}^{2}$, closed-form expressions are still available in the following cases.

### v2

@@ IV-C Analytical Optimization under Weak Uniform Fields
If the external magnetic field remains unchanged, the deformation of MeSCRs is determined by the placement (position) and orientation of embedded magnets. For instance, embedded magnets with variable positions have been employed to expand the MeSCR’s workspace [30]. In the PRB model, the length of embedded magnets is defined by the Voronoi length as $L_{k}\triangleq\sum_{i\leq k}\ell_{i},\forall k\in\mathbb{K}$. From (17), it follows that the flexible segment beyond the last embedded magnet is unactuated. It does not affect the workspace volume of the MeSCR, nor any performance metric defined using a left-invariant Riemannian metric. Therefore, we let one of the magnets be fixed at the distal end $(L_{k_{N_{m}-1}}=L)$, leaving $N_{m}-1$ decision variables on placements. Additionally, we let the flexible segments between adjacent magnets have minimal gaps $(L_{g})$ to mitigate magnetic coupling. The orientation of embedded magnets $(\bar{\mathbf{m}}_{k}\in\mathbb{S}^{2})$ is also considered as the design variable.
We collect the design variables of magnet placements and orientations into the parameter vector $\mathbf{a}\in\mathbb{A}$. Under a spatially uniform magnetic field $\mathbf{b}\in\mathbb{B}$ serving as the actuation source, the structural optimization of the MeSCR follows the general form of (9), where the equilibrium mapping $\boldsymbol{\theta}^{\ast}(\mathbf{b};\mathbf{a})$ is implicitly defined by the fixed-point relation in (17). Suppose the magnetic field $\mathbf{b}$ satisfies the condition in Theorem 1. The manifold $\mathbb{B}$ then constitutes an affine subspace diffeomorphic to $\mathbb{R}^{3}$. By Theorem 1, the stiffness matrix $\mathbf{S}$ is nonsingular in a neighborhood of each $\mathbf{b}_{u}\in\mathbb{B}$, ensuring that the equilibrium mapping is differentiable with respect to $\mathbf{b}$. Furthermore, by Theorem 2, the Jacobian $\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}}=\mathbf{S}^{-1}\mathbf{M}$ satisfies $\operatorname{rk}(\mathbf{S}^{-1}\mathbf{M})=2N_{m}$, implying that $\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}_{u}}=\mathbf{S}^{-1}\mathbf{M}\mathbf{U}_{u}$ is of full rank when $N_{m}\geq 2$. Consequently, the immersion $\boldsymbol{\theta}^{\ast}:\mathbb{B}\to\mathbb{T}$ is a smooth mapping, and the structural optimization of the MeSCR can be expressed in the analytical integral form of (12) derived under the smooth manifold assumption.
@@
Since $\mathcal{J}_{\boldsymbol{\theta}}$ contains complex Hessian terms, we first derive a tractable approximation to facilitate further analysis.
@@
$\mathcal{J}_{\boldsymbol{\theta}}^{2}=\det\left(\mathbf{U}_{u}^{\top}\mathbf{M}^{\top}\boldsymbol{\Lambda}^{-2}\mathbf{M}\mathbf{U}_{u}\right)+\epsilon\triangleq\tilde{\mathcal{J}}_{\boldsymbol{\theta}}^{2}+\epsilon$ (21)
@@
It is deferred to the supplementary material. ∎
It shows that $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}$ provides a close approximation of $\mathcal{J}_{\boldsymbol{\theta}}$ in the weak-field regime, with the residual term converging to zero as the magnetic field strength decreases. We can now turn to analyze the approximation of the objective $\tilde{\mathcal{Z}}(\mathbf{a})\triangleq\int_{\mathbb{B}}z\ \tilde{\mathcal{J}}_{\boldsymbol{\theta}}\ \Omega_{\mathbb{B}}$. Let $\mathbf{M}(\boldsymbol{\theta})$ in (S15) be represented as $\mathbf{M}=\mathbf{D}_{k_{m}-1}^{\top}\mathbf{U}_{m}\mathbf{D}_{m}$, we notice
@@
with $\Lambda_{k_{i}}^{m}=\sum_{j=0}^{k_{i}}\frac{\ell_{j}^{2}}{E_{j}^{2}I_{j}^{2}}\approx\frac{1}{k_{0}E^{2}I^{2}}\sum_{j=0}^{i}(L_{k_{j}}-L_{k_{j-1}})^{2}$ as a function of $L_{k}$. Lemma 1 also derives the approximation:
$=\sum_{ij}\Lambda_{ij}^{m}\left(\mathbf{m}_{k_{i}}\mathbf{m}_{k_{j}}^{\top}-\left(\mathbf{m}_{k_{i}}^{\top}\mathbf{m}_{k_{j}}\right)\mathbf{I}_{3}\right).$
The following lemma shows the relation between the singularity of $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}$ and the configuration of $L_{k}$:
@@
It is deferred to the supplementary material. ∎
@@
If the performance index satisfies $z(\mathbf{b}_{u};L_{k})>0$, and the structural parameter $L_{k}$ satisfies the boundary configuration condition in Lemma 2, then the objective $\tilde{\mathcal{Z}}$ attains its minimum at this configuration.
@@
Since $z>0$ and $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}\geq 0$, it follows that $\tilde{\mathcal{Z}}\geq 0$ for any admissible $L_{k}$. When $L_{k}$ satisfies the boundary configuration condition in Lemma 4, we have $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}(\mathbf{b})\equiv 0$ over $\mathbb{B}$, and thus $\tilde{\mathcal{Z}}=0$. ∎
@@
Due to the Jacobian singularity at boundary configurations, the value of the objective becomes insensitive to the gradient of the performance index. However, the minimum spacing constraint prevents the structural parameter $L_{k}$ from reaching such degenerate configurations. Since $\mathcal{J}_{\boldsymbol{\theta}}$ remains continuous on $L_{k}$, it stays small in the vicinity of the boundary, making nearby configurations highly competitive in minimizing the objective. In this case, the performance gradient becomes the key factor in selecting specific configurations. One can also reduce the optimization bias by considering the objective as $\mathcal{Z}_{\mathrm{vol}}/\mathcal{Z}_{\mathrm{dex}}$. This normalized objective reflects the average performance on the configuration space and is insensitive to Jacobian singularity.
Example: Consider a MeSCR with two axially embedded magnets actuated by a spatially uniform field $\mathbf{b}_{u}$ with $\mathbf{s}^{\top}\bar{\mathbf{t}}=0$, where the global manipulability is to be maximized [21]. The embedded magnets have identical dipole strength. Let one magnet be fixed at the distal end ($L_{k_{1}}=L$), while the position of the other magnet $L_{k_{0}}$ serves as the design variable, subject to the boundary constraint $0<L_{k_{0}}<L$. For planar deflection, the joint variable simplifies to $\boldsymbol{\theta}_{i}=\theta_{i}\mathbf{n}$, with $\mathbf{n}\in\mathbb{S}^{2}$ represents the normal vector. In this configuration, the MeSCR becomes kinematically equivalent to a planar revolute open-chain mechanism. Let $\theta_{i}^{j}\triangleq\sum_{l=i}^{j}\theta_{l}$ denote the cumulative joint angle. In the regime of weak magnetic actuation, the global manipulability index has the linearized expression on $\theta_{k_{0}}^{k_{1}}$: $z\approx L_{k_{0}}(L-L_{k_{0}})\left|\sin\theta_{k_{0}}^{k_{1}}\right|$, where the analytical form is derived from the planar 2-joint revolute open-chain mechanism.
IV-C1 Optimal Orientation
We parameterize the in-plane magnetic moments by angles $\{\varphi_{k}\}$ with respect to the $\bar{\mathbf{v}}$-axis, where the initial orientation is $\bar{\mathbf{t}}$. Let $\varphi_{k_{0}}^{k_{1}}\!\triangleq\!\varphi_{k_{1}}-\varphi_{k_{0}}$ denote the relative orientation, and assume a spatially uniform planar field $\mathbf{b}_{u}\!\in\!\mathbb{R}^{3}$ with support on the $\bar{\mathbf{t}}$-$\bar{\mathbf{u}}$ plane. Under the weak-field regime and planar bending, the immersion Jacobian admits the linearized form $\mathcal{J}_{\boldsymbol{\theta}}\;\approx\;c_{0}\big|\sin\!\big(\varphi_{k_{0}}^{k_{1}}+\theta_{k_{0}}^{k_{1}}\big)\big|,$ where $c_{0}>0$ collects geometry- and stiffness-related constants. The corresponding global performance functional can be written as
$\mathcal{Z}\!\left(\varphi_{k_{0}}^{k_{1}}\right)\;=\;c_{1}\!(\sin(\varphi_{k_{0}}^{k_{1}}+2\theta_{k_{0}}^{k_{1}})/4\;-\;\theta_{k_{0}}^{k_{1}}\cos(\varphi_{k_{0}}^{k_{1}})/2),$ (23)
with $c_{1}>0$ independent of $\varphi_{k_{0}}^{k_{1}}$. In the small-angle limit, the stationarity condition leads to
$\lim_{\theta_{k_{0}}^{k_{1}}\to 0}\nabla_{\varphi_{k_{0}}^{k_{1}}}\mathcal{Z}=0\;\;\Longrightarrow\;\;\big(\varphi_{k_{0}}^{k_{1}}\big)^{\ast}\;=\;\frac{\pi}{2}\;(\text{mod }\pi),$ (24)
indicating that the optimal in-plane orientation is orthogonal in the uniform weak field regime. Interestingly, it is analogous to the orthogonal joint arrangement in serial manipulators, where perpendicular joint axes maximize manipulability by decoupling motion directions around the nominal configuration. Similar orthogonal magnetization has been widely adopted in radially magnetized catheter robots [31, 32], where the magnetic moment is aligned perpendicular to the backbone to enhance steering torque and directional responsiveness under uniform magnetic fields.
IV-C2 Optimal Placement
We focus on axially magnetized configurations, as they are the most common design adopted in continuum magnetic robots. Applying the approximation from (22), the immersion Jacobian also admits the linearized form:
$\mathcal{J}_{\boldsymbol{\theta}}\approx$
[…]
The cumulative rotation angle can also be approximated using (20) as $\left|\theta_{k_{0}}^{k_{1}}\right|\approx\frac{L-L_{k_{0}}}{EI}\left|\bar{\mathbf{m}}_{k_{1}}^{\top}[\mathbf{n}]_{\times}^{\top}\mathbf{b}_{u}\right|$. Substituting this into the above expressions for $z$ and $\mathcal{J}_{\boldsymbol{\theta}}$, we can factor the dependence on $L_{k_{0}}$ out of the integral in the objective. The structural optimization problem thus reduces to:
[…]
$L_{k_{0}}^{\ast}\approx\begin{cases}1/2.718\ L,\ \text{if }\bar{\mathbf{m}}_{k_{0}}=\bar{\mathbf{m}}_{k_{1}},\\ 2/7\ L,\ \text{if }\bar{\mathbf{m}}_{k_{0}}=-\bar{\mathbf{m}}_{k_{1}}.\end{cases}$